Normal Atmospheric Pressure is 101, 300 Pa. Oil fl
Answer #1

P1 = Patm = 101300 Pa

----------

P2 = P1 + rho g H

P2 = (101300) + (850 x 9.81 x 11) = 193023.5 Pa


------------

P3 = P2 + rho g h

P3 = (193023.5) + (1030 x 9.81 x 10) = 294066.5 Pa